If X = 27 inches, Y = 120 inches, and Z = 123 inches, what is the cosine of angle B?

Answer by Edwin McCravy(20060)

We check to see if it is a right triangle by
the Pythagorean theorem.

    X² + Y² ?=? Z²

 27² + 120² ?=? 123²

729 + 14400 ?=? 15129

      15129  =  15129

So yes it is a right triangle.  Therefore,

longest side = hypotenuse = Z = 123 inches
side opposite A = X = side adjacent angle B = 27 inches
side opposite B = Y = side adjacent angle A = 120 inches

  

Edwin